At a discount of 3% for a one- year treasury bill, an investor paid 970.87 pesos. How much will the investor received at maturity??
1
Expert's answer
2020-11-15T18:35:49-0500

"p=970.87(1-\frac{dr}{360})"


"=970.87(1-\frac{0.03}{360})"


"=970.87(1-0.00008333)"


"970.87\times0.99991667"


"=970.7891"

Because you're buying a 970.87 pesos Treasury bill instead of one for 100 pesos, multiply 970.7891 by 10 to get the final price of 9707.891 pesos